Could someone please show me how to find an equation of the line containing the given pair points (-2,-7) and (-9,-3) The equation of the line in slope intercept from y__
Answer by stanbon(75887)

Could someone please show me how to find an equation of the line containing the given pair points (-2,-7) and (-9,-3) The equation of the line in slope intercept from y__
--------------
slope = (-3--7)/-9--2) = 4/-7 = -4/7
-------------------
You have y = -7 when x = -3 and m = -4/7
---
Form: y = mx + b
Substitute and solve for "b":
-7 = (-4/7)(-3) + b
-7 = (12/7) + b
---
-49/7 = 12/7 + b
b = -61/7
=============
Cheers,
Stan H.
=============